### v2

# AVA-Bench: Atomic Visual Ability Benchmark for Vision Foundation Models

Abstract
[…]
Figure 1: Vision foundation models (VFMs) trained with different data and objectives are evaluated on the proposed AVA-Bench to assess their strengths and limitations across atomic visual abilities (AVAs).
1 Introduction
Figure 2: Visual Question Answering (VQA) often requires multiple atomic visual abilities (AVAs) to answer a question. As such, when a model makes an incorrect prediction, it can be difficult to determine whether it stems from a failure to capture all required AVAs or just a single critical one.
Vision Foundation Models (VFMs), pre-trained on large and diverse datasets, have become central to AI by providing transferable features for a wide range of downstream tasks khan2022transformers ; awais2025foundation ; chowdhury2025prompt ; bommasani2021opportunities . The variety of pre-training objectives and supervision signals has led to a proliferation of specialized VFMs—such as DINOv2 oquab2023dinov2 , CLIP radford2021learning , and SAM kirillov2023segment —each excelling in distinct visual capabilities while often exhibiting interesting emergent properties goldblum2023battle ; caron2021emerging ; naseer2021intriguing . Consequently, establishing a systematic and effective evaluation protocol for VFMs has become increasingly crucial.
Existing evaluation protocols can generally be categorized into two groups. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification han2022survey ; mai2024fine and COCO for detection or segmentation thisanke2023semantic ; balachandran2024eureka . To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks liu2023visual ; zhu2023minigpt ; chowdhery2023palm .
While increasingly adopted, this generic protocol suffers from two potential blind spots: (i) discrepancies between the instruction tuning data and test VQA data can cause performance drops due to data mismatch rather than genuine visual limitations in VFMs; and (ii) existing VQA benchmarks typically require multiple visual abilities simultaneously, making it difficult to determine whether a failure arises from the absence of multiple abilities or merely a single critical one.
To address these challenges, we introduce AVA-Bench, the first VFM evaluation benchmark explicitly designed to disentangle Atomic Visual Abilities (AVAs)—the fundamental visual capabilities that combine to solve complex visual reasoning tasks. For instance, answering typical VQA questions shown in Figure 2 necessitates integrating several AVAs.
Specifically, AVA-Bench evaluates VFMs across 14 carefully identified AVAs (see Figure 3), including localization, counting, spatial reasoning, orientation, absolute and relative depth estimation, and recognition of textures, colors, objects, actions, emotions, optical characters (OCR), and scenes. Each AVA comes with distribution-matched train and test splits and is probed in isolation, eliminating the two aforementioned ambiguities. This enables AVA-Bench to pinpoint exactly where a VFM excels or falters, providing a clear picture of its strengths and weaknesses.
We systematically benchmark leading VFMs trained under diverse objectives and data (see Figure 1), covering language-supervised (e.g., SigLIP-1/2 tschannen2025siglip ; zhai2023sigmoid , CLIP radford2021learning , InternVL-2.5 chen2024expanding ), multimodal autoregressive (AIMv2 fini2024multimodal ), segmentation-supervised (SAM kirillov2023segment ), depth-supervised (MiDas ranftl2020towards ), contrastive self-supervised (DINOv2 oquab2023dinov2 ), and agglomerative models (RADIO ranzinger2024radio ). We follow the standard protocol of adding an LLM on top, but fine-tune it separately for each AVA.
Our extensive analyses reveal the following main findings: (1) SigLIP-1/2 and AIMv2 emerge as the most versatile VFMs, achieving the highest average rank across all AVAs—highlighting the critical role of language supervision in enhancing general visual capability; (2) For vision-centric AVAs such as localization, absolute depth estimation, and orientation, the SSL-based DINOv2 performs comparably or better than language-supervised counterparts; (3) Conversely, language-centric tasks such as OCR strongly favor language-supervised VFMs; (4) Last but not least, we observe that VFMs universally excel at low- to mid-level AVAs (e.g., texture, relative depth estimation, object recognition), regardless of their training objectives—suggesting that VQA failures typically stem from deficiencies in specific critical AVAs rather than broad visual incompetence. These insights turn VFM selection from educated guesswork into principled engineering, enabling practitioners to choose (or ensemble) VFMs based on the specific AVA strengths required by their downstream tasks.
[…]
Contributions. Our key contributions are three-fold:
  • We identify critical blind spots in existing evaluation protocols and introduce AVA-Bench, a systematic, diagnostic, and comprehensive VFM evaluation benchmark covering 14 atomic visual abilities (AVAs), clearly highlighting VFMs’ fundamental strengths and weaknesses.
  • We conduct a detailed evaluation and insightful analysis of diverse leading VFMs, deriving actionable guidance for VFM selection in downstream applications such as customized MLLMs.
  • We release a resource-efficient evaluation protocol along with an open-source codebase to facilitate the development of the next generation of accountable and versatile VFMs.
2 Vision Foundation Models (VFMs) and Evaluation
Figure 3: AVA-Bench consists of 14 Atomic Visual Abilities (AVAs) that can be combined to address more complex visual reasoning tasks.
2.1 Vision Foundation Models (VFMs)
A wide range of architectures and learning strategies have been explored to develop vision foundation models (VFMs). In general, they fall into two categories based on the nature of their training data. On one hand, VFMs are pre-trained purely on visual data. For instance, the Vision Transformer (ViT)  dosovitskiy2020image is trained on labeled images using the supervised signal to capture visual representations. DINOv2  oquab2023dinov2 , in contrast, adopts a self-supervised learning approach, enabling the model to learn those features without labeling. Some VFMs are designed for specific vision tasks—for example, SAM kirillov2023segment specializes in open-vocabulary segmentation, while MiDaS ranftl2020towards focuses on monocular depth estimation. RADIO  ranzinger2024radio instead introduces a multi-teacher distillation framework that unifies the strengths of different VFMs (e.g., CLIP, DINOv2, and SAM) into a single efficient student model.
Another family of VFMs leverages image-text pairs. CLIP  radford2021learning uses contrastive learning to align image and textual descriptions, while SigLIP  tschannen2025siglip ; zhai2023sigmoid replaces the contrastive loss with a sigmoid one for more efficient training. Unlike the conventional language-guided VFMs that process images and text separately, AIMv2  fini2024multimodal integrates visual and textual understanding into a single auto-regressive framework, which is simple yet effective. Due to their different architectures and learning objectives, VFMs may have varying strengths and limitations in visual understanding.
2.2 LLM-based Evaluation for VFMs
Figure 4: Evaluation pipeline for AVA-Bench: The standard LLaVA-style two-stage training prepares the connector and LLM for VFM evaluation. For each AVA, only connector and LoRA is trained.
Figure 5: Overall statistics of AVA-Bench.
Unlike the traditional paradigm, where different perception tasks, such as classification, detection or segmentation, necessitate task-specific models han2022survey ; awais2025foundation , the widespread adoption of Large Language Models (LLMs) as versatile interfaces has significantly shifted this paradigm zhang2024mm ; wu2023multimodal . Reflecting this shift, recent studies advocate leveraging large language models (LLMs) as general-purpose heads and evaluating VFMs on broad Visual Question Answering (VQA) benchmarks tong2024cambrian . Specifically, following the LLaVA approach, these studies utilize a two-stage framework: (i) pre-training a connector between frozen LLM and VFMs with image-text pairs for feature alignment (ii) fine-tuning both the connector and the LLM using instruction-tuning data, while keeping VFMs frozen during both stages (see Figure 5). This LLM-based evaluation protocol has rapidly gained popularity as it closely mirrors the contemporary multimodal LLM setting and effectively captures diverse real-world perception challenges.
However, this evaluation protocol suffers from two potential blind spots. First, discrepancies may exist between instruction-tuning datasets and the test VQA datasets. Consequently, a miss-prediction might arise from data mismatch rather than genuine visual deficiencies in VFMs. Second, typical VQA benchmarks typically require multiple visual abilities simultaneously to produce correct answers. This multi-ability requirement obscures the exact cause behind a model’s incorrect predictions. For instance, if a VFM is proficient in almost all visual abilities except orientation recognition (forward or backward), it would fail on the question in Figure 2 (left). In such scenarios, current evaluation methods only provide relative comparisons between VFMs without elucidating the specific missing visual capabilities. Thus, our AVA-Bench aims to complement the popular LLM-based evaluation by pinpointing exactly where a VFM excels or falters, yielding a holistic understanding of their strengths and shortcomings.
3 AVA-Bench
Figure 6: Performance comparison of VFMs across all AVAs. (Left) Language-Supervised VFMs with DINOv2 as a reference. (Right) Other VFMs with the SigLIP-2 as a reference.
To address the blind spots discussed earlier, we introduce AVA-Bench, the first systematic evaluation suite explicitly designed to disentangle 14 fundamental perceptual skills–Atomic Visual Abilities (AVAs)–for VFMs. We first introduce the concept of AVAs (subsection 3.1), followed by detailing our dataset curation pipeline (subsection 3.2). Lastly, we outline the overall dataset statistics and discuss quality control (subsection 3.3).
3.1 Atomic Visual Abilities (AVAs)
AVAs are fundamental perceptual capabilities that can be combined to address more complex visual reasoning tasks. Rather than treating VQA as a monolithic task, we break it down into 14 fundamental perceptual abilities–such as counting, depth estimation, localization and spatial reasoning–each of which can be composed to answer more complex questions. For example, answering the question, “How many yellow dogs are facing backward on the left-hand side of the stop sign?”, requires multiple AVAs: counting, color recognition, localization, and spatial reasoning (see Figure 2). By explicitly defining the set of AVAs required to interpret a question, our benchmark quantitatively characterizes where a VFM excels or falters across each core ability.
The AVAs selected for AVA-Bench are grounded in a thorough literature analysis, including compositional text-to-images benchmarks huang2023t2i ; kil2024mllm ; wu2024conceptmix and VQA questions goyal2017making ; RealWorldQA ; ainslie2023gqa (details in Appendix), focusing strictly on pure perceptual tasks and excluding non-perceptual reasoning skills (e.g., historical context, mathematical reasoning). Examples of each AVA can be found in Figure 3 and detailed definitions can be found in Appendix.
3.2 Dataset Curation
Constructing AVA-Bench required carefully isolating image–question pairs that specifically test individual AVAs. Existing MLLM and VQA datasets often blend multiple perceptual abilities, complicating direct assessments tong2024eyes ; zhang2024mme ; yu2023mm . To achieve clear isolation, we assembled a comprehensive suite comprising image–question pairs from 26 diverse datasets, explicitly aligning each pair with a single targeted AVA. These datasets span a broad range of domains—including general scenes, wildlife (e.g., birds, fungi, plants), vehicles, indoor/outdoor settings, and remote-sensing imagery.
To construct the benchmark for an AVA, image–question pairs were carefully designed or adapted to focus solely on that AVA. For instance, the original question “What’s the depth of the car from the camera?” involves both localization and depth estimation. To isolate depth estimation, we explicitly provide the car’s bounding box. This approach enables fine-grained diagnostic insights into VFMs’ AVA-specific strengths and weaknesses. Detailed examples are illustrated in Figure 3, with further details specific below and in the Appendix.
3.2.1 Localization
Localization is a crucial foundational AVA, essential for various VQA questions. We curate 34.8K localization-focused image–question pairs sourced from diverse datasets: Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote-sensing). To ensure clarity, we include images with a single instance of the target object and exclude objects with bounding box area smaller than 0.2% of the image.
3.2.2 Counting
We collect 11.3K image-question pairs from five datasets: VQAv2 goyal2017making (open-domain), FSC ranjan2021learning (open-domain counting), CARPK hsieh2017drone (cars), Crowd Surveillance Dataset li2022video (people), and LVIS gupta2019lvis (open-domain). FSC, CARPK, and Crowd datasets explicitly cater to counting tasks, ensuring clear and distinct instance counts. LVIS instance masks are utilized to generate additional counting questions, complemented by counting-focused selections from VQAv2.
3.2.3 Spatial Reasoning
We curate 15.9K pairs from NYU-Depth V2 silberman2012indoor for indoor scenes with object segmentation, and LVIS gupta2019lvis and Objects365 shao2019objects365 for diverse open-domain scenarios. For each image, we select two non-overlapping objects—one marked with a blue bounding box (the reference object) and another with a red bounding box (the target). The model is then asked to determine the relative spatial position of the target object with respect to the reference one, choosing from multiple-choice options: top-left, top-right, bottom-left, and bottom-right. We ensure that the two boxes do not overlap either horizontally or vertically to avoid ambiguity.
3.2.4 Visual Attribute
Orientation. We compile 10.1K pairs from two specialized datasets, CURE-OR Temel2018_ICMLA and EgoOrientBench jung2024right , providing uncluttered images of objects from nine distinct orientations: front, back, left, right, top, front left, front right, back left, and back right.
Color. We curate 14K pair sourcing from from open-domain datasets, such as Objects365 shao2019objects365 and LVIS gupta2019lvis . To isolate the color recognition from object semantics or background clutter, we provide bounding boxes that tightly localize regions with minimal color variance, ensuring that the model focuses on RGB prediction.
Texture. We collect 13.2K pairs from four open-domain datasets: DTD cimpoi14describing , Kylberg kylberg2011kylberg , KTH-TIPS kth-tips_dataset and KTH-TIPS2 article with diverse texture types, such as striped, aluminum foil, and zigzagged.
Emotion. We gather 13.9K pairs from two publicly available datasets, RAF-DB li2019reliable and ExpW lian2020expression , covering 7 annotated human emotional states: surprise, neutral, disgust, fear, happy, sad, and angry.
3.2.5 Depth Estimation
We curate depth-focused image-question pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes.
Absolute Depth. We assemble 9K pairs and for each sample, we place a bounding box on a target object and ask the model to estimate its distance from the camera. We ensure that each object class spans at least five distinct depth bins and that the samples within each bin are balanced.
Relative Depth. We collect 11.5K pairs. Each image is annotated with two non-overlapping bounding boxes for two distinct objects. The model is asked to determine which object is closer to the camera. To prevent annotation bias, we ensure that each object class is evenly distributed between being the nearer or farther object to avoid cases where certain objects are always closer.
3.2.6 Recognition
Action. We construct 15K pairs from the Moments in Time monfort2019moments , a video dataset with diverse human actions. We extract the middle frame from each 3-second video for 302 distinct actions
Fine-grained. We curate 9K pairs from three diverse datasets: CUB-200-2011 wah2011caltech (birds), iNaturalist-2021 van2021benchmarking (fungi, plants, animals), and FGVC-Aircraft maji2013fine (objects). We randomly select 50 species across fungi, plants, and animals from iNaturalist van2021benchmarking and 100 bird species from CUB, resulting in a total of 300 fine-grained classes.
Object. We curate 34.8K pairs from 4 datasets with diverse domains across 70 unique objects: Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote sensing). For each image, we provide a bounding box to eliminate the need for localization.
Scene. We curate 13.9K pairs from two datasets: Places434 zhou2017places (open-domain) and AID xia2017aid (remote sensing). The model is required to select the correct scene class from 30 randomly sampled options. The full set includes 464 classes spanning a wide range of scenes.
Figure 7: The ranks of VFMs across all AVAs. Detailed results in Appendix.
3.3 Quality Control and Dataset Statistics
We emphasize rigorous quality control to ensure fair, balanced, and unbiased assessments. Every AVA follows an 80/20 split in which the exact object classes and answer bins that appear in training are mirrored in testing. This ensures that performance differences truly reflect the VFM’s perceptual capabilities rather than train-test distribution mismatches, effectively addressing the concerns highlighted in subsection 2.2. Moreover, we try to avoid potential bias in the benchmark. For example, in counting, we explicitly balance the number of samples per counting bin and per object type in both training and testing sets. This approach avoids biases wherein certain counts (e.g., “7 apples”) might dominate the training data, artificially inflating accuracy for specific numerical predictions. In another example, for localization, we set a clear threshold for minimum bounding-box area to ensure object visibility. More details about how we ensure the quality of AVA-Bench can be found in Appendix.
To further improve the generalizability within each AVA, we aggregate data from multiple diverse datasets. This ensures that models encounter the same core visual ability across varied scenes, object types, and data distributions, making the assessment more robust and less dataset-specific. Lastly, image-question pairs are intentionally crafted or selected to be simple, clear, and explicitly focused on testing only one AVA at a time. In summary, AVA-Bench comprises 218K meticulously curated image-question pairs that robustly isolate individual AVAs, carefully control dataset balance and visibility, and systematically prevent annotation biases. Detailed statistics are provided in Figure 5, with additional details in Appendix.
4 Evaluation Pipeline of AVA-Bench
To evaluate VFMs using AVA-Bench, we adopt the established LLM-based VFM evaluation protocol, employing the standard LLaVA-style two-stage training procedure to prepare the connector and LLM for VFM evaluation (details in subsection 2.2 and Figure 5). For each AVA in AVA-Bench, we fine-tune the connector and LLM while keeping the VFM frozen. Given the modest size of the training sets per AVA (typically around 6K–10K), we employ Parameter-Efficient Fine-Tuning (PEFT) mai2025lessons ; houlsby2019parameter ; tu2023visual , specifically Low-Rank Adaptation (LoRA) hu2022lora , to mitigate potential overfitting. Subsequently, the fine-tuned model is evaluated on the corresponding AVA-specific test sets.
4.1 Is a Heavyweight LLM Evaluator Necessary?
[…]
5 Experiments
5.1 Setup
All experiment setup follow the evaluation pipeline described in section 4. More experiment details can be found in Appendix.
Metrics for AVAs. For absolute depth and counting AVAs, we utilized a normalized mean absolute error (MAE) relative to the ground-truth. This normalization ensures that errors involving greater distances or counts, which are inherently more challenging, are proportionally penalized less severely. Localization performance is evaluated using Generalized Intersection-over-Union (GIoU rezatofighi2019generalized ), color via CIEDE2000 luo2001development , and OCR through Average Normalized Levenshtein Similarity (ANLS) biten2019scene . All other AVAs employ standard accuracy metrics. Please refer to Appendix for more metric details.
5.2 Observations and Analyses
Figure 9: Impact of bounding boxes on spatial reasoning performance. With bounding boxes, all VFMs perform perfectly in spatial AVA; without them, models with weaker localization (MiDaS, SAM) perform worse.
Bounding boxes help isolate and assess specific AVAs. Bounding boxes serve as a crucial tool for disentangling AVAs, allowing us to isolate and evaluate specific capabilities rather than compounded performance on complex tasks. For example, when assessing spatial reasoning, VFMs exhibit uniformly strong and similar performance when provided with ground-truth bounding boxes, indicating a shared understanding of object relationships. However, removing the bounding box input transforms this into a composite task requiring both localization and spatial reasoning. As shown in Figure 9, this removal leads to a significant divergence in model performance. While SigLIP-2 and DINOv2 maintain high performance, Midas and SAM perform considerably worse in this composite setting. Analyzing their localization performance independently reveals that Midas and SAM possess significantly weaker localization capabilities compared to SigLIP-2 and DINOv2. This stark contrast confirms that performance degradation on composite tasks is often attributable to deficiencies in specific, untrained or underdeveloped constituent AVAs, such as localization in the case of Midas and SAM, rather than a fundamental failure in all visual understanding capacities.
Subgroup analyses reveal exceptions hidden by aggregate metrics. Aggregate metrics can sometimes obscure important nuanced insights. To gain a deeper understanding, we conduct detailed analyses by partitioning test samples based on specific criteria (e.g., object size in localization tasks) and examining whether these subgroup trends align with overall performance. We split localization testing samples based on normalized bounding box sizes (relative to image size), where 0.1 indicates an object occupies 10% of the image area. As illustrated in Figure 10, VFMs surprisingly exhibit minimal performance differences when localizing large objects (0.3–0.5). Conversely, performance disparities amplify as object size decreases, revealing significant weaknesses in MiDas and SAM for smaller objects. More subgroup analyses for other AVAs can be found in the Appendix.
Niche mastery of a modest model. Interestingly, almost all VFMs, including those with generally lower performance, excel in at least one specific AVA (Figure 7). For example, SAM achieves exceptional results in color recognition, and DINOv2 excels notably in orientation.
Consistently good VFM performance across lower and mid-level AVAs. All VFMs, regardless of their training strategies, demonstrate good performance in low- to mid-level AVAs such as texture recognition, relative depth estimation, and object recognition ( Figure 6). This uniformity implies that failures in complex visual reasoning predominantly arise from deficiencies in specific, critical AVAs rather than general shortcomings in visual understanding.
Language-aligned pretraining boosts performance on language-centric AVAs. AVAs that involve understanding visual information intertwined with text, such as OCR, substantially benefit from language-aligned pretraining. Non-language-aligned VFMs significantly underperform in these tasks, as illustrated in Figure 6, highlighting the importance of language alignment in VFM training.
The role of language supervision in VFM versatility. Language-supervised VFMs, specifically SigLIP-1/2 and AIMv2, demonstrate broad competency across AVAs (Figure 7 & Figure 6). Their consistent high ranking highlights that language supervision is crucial to developing general-purpose visual reasoning abilities.
Figure 10: Detail results for localization for overall and different splits based on the ground-truth’s bounding box size normalized by image size. For example, 0.1 means the bounding box size is 10% of the image size. Higher GIoU is better.
6 Discussion
Struggles of Non-Language-Aligned VFMs. Figure 7 and Figure 6 indicate that non-language-aligned VFMs, despite exhibiting strengths in certain vision-centric or low-to-mid-level AVAs (e.g., DINOv2 excelling in orientation and SAM leadning in color), typically underperform in most AVAs. To investigate the underlying cause of these shortcomings, we conducted a preliminary analysis focusing on the visual feature representations before and after the LLM connector. Using linear probing on max-pooled visual features from DINOv2, we observed a substantial accuracy degradation from 66.3% (pre-connector) to 25.67% (post-connector) for the fine-grained recognition AVA, aligning with findings reported in previous work finer . This sharp performance drop suggests that critical visual information is often compromised during modality alignment processes. Recent studies have shown that partially unfreezing and fine-tuning the last few layers of VFMs can enhance performance chen2024sharegpt4v ; however, this approach risks eroding the generalizability that initially makes these VFMs valuable. Alternatively, agglomerative models ranzinger2024radio ; heinrich2024radio such as RADIO-2.1 demonstrate relatively robust performance on AVA-Bench, suggesting potential in combining specialized VFMs. Nevertheless, effectively aligning non-language-aligned VFMs to language modalities without sacrificing their inherent visual strengths remains a challenge, highlighting an important avenue for future research.
Platonic Representation Hypothesis Holds? Recent research suggests that large-scale training may lead VFMs toward converging onto similar representational spaces–Platonic Representation Hypothesis platonic . Our findings partially support this hypothesis: for certain AVAs such as object recognition, texture, and relative depth estimation, VFMs demonstrate universally strong performance, indicative of similar underlying visual representations irrespective of training objectives. Conversely, significant performance disparities among VFMs in most other AVAs, as revealed by our evaluation, indicate limitations to the generality of this hypothesis. Thus, our study suggests that the Platonic Representation Hypothesis might only hold in certain cases and warrants further empirical scrutiny in more diverse and challenging contexts.
What Are Things Going From Here? While MLLM have demonstrated remarkable versatility, they are not universally effective in all scenarios, especially in specialized domains cheng2024domain ; liang2024comprehensive . Thus, there is a growing necessity for developing specialized MLLMs kumar2024diagnostics ; li2025visual . Currently, selecting appropriate VFMs for such customized MLLMs remains largely heuristic tong2024cambrian ; sun2025multi ; li2025openvision . Our work provides actionable insights that transform this selection process from heuristic guesswork into principled engineering. By clearly identifying AVA-specific strengths and weaknesses, practitioners can now systematically choose VFMs to precisely address the particular visual demands of targeted downstream tasks. Moreover, AVA-Bench represents a critical step towards developing next-generation VFMs by providing a systematic, diagnostic, and comprehensive evaluation framework. This benchmark enables VFM developers to accurately pinpoint specific deficiencies and implement targeted improvements, fostering the creation of more robust, versatile, and well-rounded VFMs in the future.
Appendix
Appendix A AVA-Bench Details
Atomic Visual Ability Definition Example Question
Counting Determining the number of instances of an object How many apples are in the image?
Localization Identifying the location of an object in the image Provide bounding box coordinate for bicycle.
Fine-Grained Differentiating between similar sub-categories of objects What species of fungi is in the image?
OCR Reading and interpreting text visible in the image What is written in the red bounding box in the image?
Absolute Depth Estimating how far an object is from the camera From the camera’s perspective, estimate how far the closest point of the car (red box) is from the camera in real-world distance, in meters.
Relative Depth Comparing distances of two objects from the camera Which object is closer to the camera, the car (red box) or the cyclist (blue box) to the camera?
Orientation Determining the facing direction or angle of an object What is the orientation of the toy bus in the image?
Spatial Inferring layout and spatial relations Considering the relative positions of two objects in the image, where is the bicycle (red box) located with respect to the towel (blue box)?
Object Recognition Identifying objects present in the image given bounding box What is in the red bounding box in the image?
Scene Recognition Identifying the broader environment or type of setting What is the scene class of the image?
Action Recognition Determining what action is being performed Which action or activity is shown in the image?
Texture Describing surface appearance or material of objects What is the texture attribute of image?
Color Identifying colors of objects given bounding box What color is shown within the bounding box?
Emotion Recognizing emotional expressions in humans given bounding box What emotion is being shown in the image?
Table 1: Atomic Visual Abilities (AVAs). We identify 14 AVAs, serving as the foundational capabilities that can be combined to tackle complex visual reasoning tasks. For each AVA, we provide the definition and an example question in ourbench.
A.1 Atomic Visual Abilities (AVAs)
As mentioned in Section 3.1, AVAs are elemental visual capabilities that can be combined to address more complex visual reasoning tasks. The definitions and representative questions for each AVA can be found in Table 1. Additional qualitative illustrations are provided in Figure 2 and Figure 5.
The 14 AVAs selected for AVA-Bench are grounded in a thorough literature analysis.
  1. 1.
    Compositional text-to-image (T2I) benchmarks. Studies on controllable generation motivate core visual primitives—number, colour, texture, object identity, spatial relations, and more—used to construct compositional prompts huang2023t2i ; wu2024conceptmix . These primitives form an initial pool of candidate abilities.
  2. 2.
    VQA question analysis. We employ GPT-4 to summarize the visual skills demanded by VQA questions in various commonly-used datasets (VQAv2 goyal2017making , RealWorldQA RealWorldQA , GQA ainslie2023gqa , etc.), thereby enriching the pool with abilities emphasized by real-world questions.
Intersecting these two sources yields a concise yet crucial set of AVAs. Moreover, we focus strictly on pure perceptual tasks and exclude non-perceptual reasoning skills (e.g., historical context, mathematical reasoning). We provide more related work discussion in Appendix D.
A.2 Dataset Curation
Spatial Reasoning wu2025spatial ; zhang2025call . We curate 15.9K image pairs from NYU-Depth V2 silberman2012indoor (indoor scenes) and LVIS gupta2019lvis and Objects365 shao2019objects365 (open-domain scenes), all containing instance segmentation annotations. In each image, two distinct, non-overlapping objects are selected—one highlighted with a blue bounding box (reference object) and another with a red bounding box (target object). The model must identify the relative spatial position of the red box with respect to the blue box, choosing from four multiple-choice options: Left above, Left below, Right above, and Right below (Figure 11). The preprocessing steps for dataset creation are summarized below:
  • To prevent ambiguity in interpretation, we restrict each image to contain only one instance of the target and reference objects.
  • Object pairs whose bounding boxes overlapped either horizontally or vertically were excluded, ensuring unambiguous assignment to the four spatial categories.
  • Extremely small bounding boxes complicating localization were filtered out. Specifically, object instances covering at least 2% of the image area for NYU-Depth V2, and at least 0.2% for LVIS and Objects365, were retained.
  • Ambiguous object categories were excluded via a manually curated list of label IDs that often lack clear boundaries or meaningful depth interpretations from NYU-Depth V2 (e.g., wall, floor, ceiling, etc.). GPT-o1 was used to unify labels across datasets, identifying common object names conveying the same meaning, and merging these samples accordingly.
  • Each object class was ensured to appear in multiple spatial positions, with 40 samples per spatial position category. This prevents models from memorizing fixed layouts.
  • For each object class and spatial position category, an 80% train and 20% test split was ensured, for uniform distribution and fair evaluation.
  • The question for each pair: “Considering the relative positions of two objects in the image, where is the microphone (annotated by the red box) located with respect to the speaker (annotated by the blue box)? Choose from A. Left above, B. Left below, C. Right above, D. Right below.”
Figure 11: Examples of Spatial Reasoning AVA Samples.
Counting yao2025countllm ; li2024llmcount . We curate a total of 11.3K images from five datasets— VQAv2 goyal2017making , FSC-147 ranjan2021learning , CARPK hsieh2017drone , LVIS gupta2019lvis , and CrowdHuman li2022video —to evaluate object counting abilities across diverse domains, including open-domain scenes, natural objects, structured environments, and densely crowded contexts. Each image is paired with a question prompting the model to count the number of instances of a specified object category, and the model must return an integer-valued answer(Figure 12). The preprocessing steps for dataset creation are summarized below:
  • To ensure valid supervision, we filter all samples to retain only those with non-zero object counts and with object_count $\leq$ 40.
  • For each object category (object id), we require at least 4–5 distinct object count values to be represented, preventing overfitting to static object layouts.
  • For each object count, we sample a fixed range of images per object_id—between 6 and 30 depending on the dataset—to balance frequency and diversity.
  • Dataset-specific sampling rules are applied:
    • VQAv2 and LVIS: 15–30 images per object count; $\geq$5 count values per object id.
    • FSC-147: 6–12 images per object count; $\geq$4 count values per object id.
    • CARPK: 10–20 images per object count; $\geq$5 count values per object id.
    • CrowdHuman: object count capped at 40; 25–50 images per count level.
  • An 80% train / 20% test split is maintained for each object id and count level to ensure balanced distribution during evaluation.
  • The question for each image is: “How many [object] are there in the image?”, where [object] refers to the annotated target category.
Figure 12: Examples of Counting AVA Samples.
Atomic Visual Abilities (AVA) Dataset Domain # Train Samples # Test Samples
Localization Objects365 shao2019objects365 Open 27.8K 6.9K
LVIS gupta2019lvis Open
iNaturalist-2021 van2021benchmarking Bird, Animal
DIOR li2020object Remote-Sensing
Counting VQAv2 goyal2017making Open 9K 2.4K
FSC ranjan2021learning Open
CARPK hsieh2017drone Car
Crowd Surveillance li2022video People
LVIS gupta2019lvis Open
Fine-grained CUB-200-2011 wah2011caltech Bird 7.2K 1.8K
iNaturalist-2021 van2021benchmarking Fungi,Plant, Animal
FGVC-Aircraft maji2013fine Object
Absolute Depth NYU-Depth V2 silberman2012indoor Indoor Scene 6.9K 1.8K
KITTI geiger2013vision Outdoor Scene
Relative Depth NYU-Depth V2 silberman2012indoor Indoor Scene 9.2K 2.4K
KITTI geiger2013vision Outdoor Scene
OCR COCO-Text veit2016cocotextdatasetbenchmarktext Open 8.8K 2.2K
IIIT5K  MishraBMVC12 Open
TextVQA balanced_vqa_v2 Open
Orientation EgoOrientBench jung2024right Open 6.9K 1.6K
CURE-OR Temel2018_ICMLA Indoor
Object Recognition Objects365 shao2019objects365 Open 27.9K 6.9K
LVIS gupta2019lvis Open
iNaturalist-2021 van2021benchmarking Bird, Animal
DIOR li2020object Remote-Sensing
Action Recognition MiT monfort2019moments Open 12.1K 3K
Texture DTD cimpoi14describing Open 10.6K 2.7K
Kylberg kylberg2011kylberg Open
KTH-TIPS kth-tips_dataset Open
KTH-TIPS2 article Open
Spatial Reasoning Objects365 shao2019objects365 Open 13K 2.9K
LVIS gupta2019lvis Open
NYU-Depth V2 silberman2012indoor Indoor Scene
Scene Recognition Places434 zhou2017places Open 11.1K 2.8K
AID xia2017aid Remote-Sensing
Emotion RAF-DB li2019reliable Human 11.9K 5.1K
ExpW lian2020expression Human
Color Objects365 shao2019objects365 Open 11.2K 2.8K
Total - - 173.5K 45.3K
Table 2: Detailed statistics of AVA-Bench.
Fine-grained paul2023simple ; gu2025bioclip ; zhang2025finer . We curate a total of 9K images from five fine-grained recognition domains—Bird, Animal, Fungi, Plant, and Object—to assess species-level recognition capabilities. The dataset sources include CUB-200-2011 wah2011caltech for birds, iNat21 van2021benchmarking for animals, fungi, and plants, and FGVC Aircraft maji2013fine for objects. Each sample contains an image and a question prompting the model to identify the specific species or object type(Figure 13). Construction details are as follows:
  • We select 100 bird species from CUB-200-2011, and 50 random classes each from the Animal, Fungi, and Plant categories of iNat21, as well as 50 classes from FGVC Aircraft for the Object category. All random selections use a fixed random seed to ensure reproducibility.
  • This results in 300 total object ids. For each class, we uniformly sample 30 images.
  • For iNat21 entries, we format species names by retaining only the last two words of their taxonomic labels for clarity and consistency.
  • For each object class, an 80% training and 20% testing split was established to ensure balanced per-class evaluation.
  • For each multiple-choice question, the candidate list includes all 50 or 100 (Object only) class names used in that task split, ensuring consistent, closed-set evaluation.
  • The question for each image is: “What species of bird is in the image? Choose one from below: 1. Cerulean_Warbler, 2. American_Crow, …, 100. Pine_Warbler ”
Figure 13: Examples of Fine-grained recognition AVA Samples.
OCR huang2025ocr ; liu2024ocrbench . We curate 10.9K images from three OCR datasets—COCO-Text veit2016cocotextdatasetbenchmarktext , IIIT5K MishraBMVC12 , and TextVQA balanced_vqa_v2 —each containing word-level bounding box annotations. In each image, a red bounding box highlights the word to be transcribed. The model is prompted to recognize the textual content inside the box based on visual context(Figure 14). The preprocessing steps for dataset creation are summarized below:
  • For COCO-Text, we retain only word-level boxes with an area greater than 1500 pixels to ensure sufficient visual resolution.
  • For TextVQA, we apply a stricter filtering criterion by retaining only word boxes with area larger than 2000 pixels.
  • For IIIT5K, we randomly sample 2,000 images from the original dataset without applying any area-based filtering.
  • Each dataset is split into 80% training and 20% validation subsets individually, before merging the resulting splits to form the final OCR benchmark.
  • A red bounding box is rendered on each image to highlight the target word location during inference.
  • The question for each image: “What is written in the red bounding box in the image?”
Figure 14: Examples of OCR AVA Samples.
Localization wu2024visionllm ; sapkota2025object . We curate 34.8K localization-focused image–question pairs sourced from Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote-sensing). Each image contains a single object instance, and the model is prompted to identify its location by providing the bounding box coordinates(Figure 15). The preprocessing steps for dataset creation are summarized below:
  • We only retain object instances whose category appears exactly once in an image, to avoid ambiguity in localization supervision.
  • We filter out objects with extremely small or large bounding boxes, retaining only those whose normalized area falls within the range of $0.002<\text{area}<0.5$ relative to the image.
  • For Objects365 and LVIS, we manually select 20 target object categories each. If the number of valid images for a category exceeds 700, we randomly sample 700 using a fixed seed for reproducibility.
  • For iNaturalist-2021, we select 10 categories from the aves (birds) class and 10 from the mammalia (mammals) class, following the same filtering and sampling strategy. All scientific names are mapped to common names to improve interpretability and model alignment.
  • For DIOR, we follow the same filtering steps and manually select 10 target categories, sampling up to 700 images per category as needed.
  • All images are padded to square format using a consistent background color computed as the mean RGB value across multiple image processors:
    $\text{background color}=RGB(124,120,111)$
    The padding preserves content aspect ratio and ensures uniform input dimensions across models.
  • For each object category, an 80% training and 20% testing split was performed after filtering and sampling, ensuring balanced and fair evaluation.
  • The question for each image follows the format: “Provide bounding box coordinate for red-tailed hawk.” The object name is dynamically replaced depending on the image.
Figure 15: Examples of Localization AVA Samples.
Recognition fu2024blink ; zhang2025revisiting ; tu2023holistic . We curate 34.8K recognition samples from four datasets spanning diverse visual domains—Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain objects), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote sensing). These samples are derived from the same images and object instances used in the localization benchmark. However, instead of asking for bounding box prediction, the recognition task requires the model to identify the object within a visually highlighted region (Figure 16).
The preprocessing steps for dataset creation are summarized below:
  • We apply the same curation strategy as in localization: only one valid instance per image, with normalized bounding box area between $0.2\%$ and $50\%$ of the image. For each dataset, 10 or 20 object categories are manually selected.
  • From Objects365 and LVIS, we select 20 object categories each, and randomly sample up to 700 images per category (using a fixed seed for reproducibility).
  • From iNaturalist-2021, we retain 10 species from the Aves (birds) and 10 from Mammalia (mammals) branches. Scientific names are converted to common English names for accessibility. Each species contributes up to 700 images.
  • From DIOR, we select 10 object categories and apply the same filtering and sampling strategy (max 700 images per class).
  • All images are padded to square shape using a consistent background color, computed from the average mean pixel values across nine vision-language processors, to ensure uniform input dimensions.
  • For each object category, an 80% training and 20% testing split was performed after filtering and sampling, ensuring balanced and fair evaluation.
  • Unlike localization, where bounding boxes are not rendered and must be predicted, in recognition the red bounding box is explicitly drawn onto each image to guide the model’s attention.
  • The question format is: “What is in the red bounding box? Choose from the following option: 1. airport, 2. american robin, …, 70. vulpes vulpes” The 70 object categories are shared across datasets and randomly shuffled for each question instance.
Figure 16: Examples of Recognition AVA Samples.
Color wang2025harnessing ; chiu2024megacoin . We curate 14K images from Objects365 shao2019objects365 and LVIS gupta2019lvis , each contributing 7K samples. This AVA focuses on assessing color perception in natural scenes. For each image, we extract a coherent color region using the following pipeline:
  • We apply SLIC superpixel segmentation and convert the image to LAB color space.
  • Superpixels with similar color values are merged to form larger regions of consistent color.
  • Among all candidate regions, we select the top-1 region with the lowest internal color variance as the final choice.
  • A red bounding box is drawn on the selected region, and the most frequent RGB color within this region is used as the answer.
  • For each object category, an 80% training and 20% testing split was performed after filtering and sampling, ensuring balanced and fair evaluation.
  • The question format for each sample is: “What color is shown within the red bounding box?”
Action szot2024grounding ; wang2025harnessing ; xie2025efficiently . We curate a total of 15K image–question pairs from the Moments in Time monfort2019moments dataset, covering a wide range of human actions and activities. Each sample is derived from a short video clip annotated with a specific action label (Figure 17). Construction details are as follows:
  • Similar action labels are merged into a unified category for clarity.
  • For each class, we randomly sample up to 500 training videos and 100 validation videos, ensuring broad yet balanced category coverage.
  • The middle frame of each selected video is extracted and used as the image representing the associated action.
  • Since listing all classes may exceed the token limits of many vision–language models, we randomly sample 100 action options per question, ensuring the ground-truth answer is always included.
  • The question format for each sample is: “Which action or activity is shown in the image? Choose from the following option: 1. buying, 2. catching, …, 100. boxing”
Figure 17: Examples of Action AVA Samples.
Emotion yang2024emollm ; li2024eald . We curate 16K image-question pairs from two large-scale facial expression datasets: RAF-DB li2019reliable and ExpW lian2020expression . These datasets consist of human portraits labeled with one of seven basic emotions: happy, sad, angry, fear, surprise, neutral, and disgust. Each image is annotated with a bounding box localizing the face of interest.
The preprocessing steps for dataset creation are summarized below:
  • Emotion labels across datasets were unified by consolidating synonymous terms (e.g., happiness and happy, anger and angry) to ensure consistent categorization across all samples.
  • Bounding box annotations provided in the datasets were used to highlight the specific individual in multi-person scenes.
  • An 80/20 train-test split was applied independently per emotion category to maintain class balance during evaluation.
  • The question for each image is framed as: “Which of the following best describes the person’s emotion in the red box? 1. happy, 2. sad, 3. angry, 4. fear, 5. surprise, 6. neutral, 7. disgust.”
Scene dai2025humanvlm ; fan2024mllm . We curate 13.9K image-question pairs from two diverse datasets: Places434 zhou2017places (open-domain) and AID xia2017aid (remote sensing). Each image is paired with a multiple-choice question, where the model selects the correct scene category from a pool of 30 randomly sampled options. The full set includes 464 unique scene classes spanning a wide range of environments(Figure 18). The preprocessing steps for dataset creation are summarized below:
  • GPT-4o was utilized to standardize labels across datasets by converting fine-grained labels into single, unified labels. Humans carefully checked each conversion to merge the newly converted labels conveying the same meaning with existing labels, ensuring semantic clarity and fluency.
  • A uniform distribution was maintained by extracting exactly 30 images per scene category, preventing category imbalance and ensuring consistent representation across classes.
  • For each scene category within both datasets, an 80% training and 20% testing split was established, ensuring balanced and fair evaluation conditions.
  • The question for each pair: “What is the scene class of the image? Choose one from below: 1. Entrance hall, 2. Lawn, … 30. Snowy Mountain.” These 30 options were selected by randomly sampling from the complete set of scene classes within each respective dataset, maintaining diversity and preventing predictable patterns.
Figure 18: Examples of Scene AVA Samples.
Texture eppel2025shape ; gavrikov2024vision . To assess a VFM’s ability to distinguish fine-grained visual patterns, we curate 13.2K image-question pairs from diverse surface textures using close-up images from four open-domain datasets: DTD cimpoi14describing , Kylberg kylberg2011kylberg , KTH-TIPS kth-tips_dataset and KTH-TIPS2 article . These datasets encompass a diverse array of texture types—such as striped, aluminum foil, and zigzagged—capturing subtle visual patterns that are essential for accurate texture recognition. Each image is paired with a multiple-choice question, requiring the model to select the correct texture label from a set of options(Figure 19). The preprocessing steps for dataset creation are summarized below:
  • GPT-4o was used to unify labels across datasets, identifying common texture attributes conveying similar meanings. These unified labels were manually verified to retain only one representative for synonymous terms.
  • Images where textures appeared as part of larger objects in cluttered scenes or within complex real-world photographs were discarded, to ensure that textures were clearly localized and recognizable without contextual interference.
  • Each texture attribute was represented by multiple images, with a minimum of 120 and a maximum of 480 samples per attribute, ensuring diversity and preventing memorization of fixed patterns by the models.
  • For each texture attribute, an 80% train and 20% test split was ensured, reaching uniform distribution and fair evaluation.
  • The question for each pair:“What is the texture attribute of the image? Choose one from below: 1. banded, 2. blotchy, …, 47. veined.” The provided options exactly match the entire option pool from each respective dataset and were shuffled to avoid bias.
Figure 19: Examples of Texture AVA Samples.
Orientation yin2025multimodal ; jung2024right . To evaluate viewpoint understanding, we curate 10.1K image-question pairs from two specialized datasets: CURE-OR Temel2018_ICMLA and EgoOrientBench jung2024right . These datasets provide uncluttered images of objects captured from nine distinct orientations—front, back, left, right, top, front left, front right, back left, and back right—allowing models to learn pose-specific cues without requiring bounding boxes(Figure 20). The preprocessing steps for dataset creation are summarized below:
  • For EgoOrientBench, each object class was ensured to appear in multiple orientations, with at least 10 and at most 40 samples per orientation label. This encourages models to learn generalized representations rather than memorizing specific arrangements.
  • From CURE-OR, we selected object instances photographed against two different background conditions using three distinct capture devices, ensuring variation in imaging style without compromising clarity.
  • For each object, 80% of the images from each orientation were assigned to the training set, and the remaining 20% to the test set, ensuring balanced representation and fair evaluation across orientations.
  • The question for each pair: "What is the orientation of the toy plane in the image? Choose one from below: 1. front, 2. front right, 3. right, 4. back right, 5. back, 6. back left, 7. left, 8. front left, 9. top". These nine options represent the common orientation labels provided by both datasets.
Figure 20: Examples of Orientation AVA Samples.
Absolute Depth xia2024large ; mi2024hierarchical ; zhang2025vision . We curate 9K image-object pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes. These datasets contain aligned RGB and depth information. Each image includes an object annotated with a bounding box, and the task requires estimating the absolute depth of the object in meters. This value is then matched against discretized ground-truth bins for evaluation(Figure 21).
The preprocessing steps for indoor dataset creation are summarized below:
  • Ambiguous object categories were excluded via a manually curated list of label IDs that often lack clear boundaries or meaningful depth interpretations (e.g., wall, floor, ceiling, etc.).
  • Images were resized to a fixed resolution of $384\times 384$, padding vertically as needed to preserve the original aspect ratio.
  • A minimum bounding box area threshold was enforced after resizing all images. Specifically, we filtered out bounding boxes smaller than 500 pixels to ensure sufficient spatial resolution for the model.
  • To ensure depth variation and avoid trivial samples, only object classes with at least 3 distinct depth bins (i.e., meaningful distribution over depth) were retained.
  • For each label, depth bins were required to have a minimum of 10 and a maximum of 30 image samples. We removed bins with insufficient data to meet this requirement and capped those with excess samples by sorting instances based on bounding box area.
  • After filtering, we retained 45 object classes, resulting in 4.4K unique image-object pairs. The dataset was split into 80% train and 20% test, preserving label and bin balance.
  • The question for each pair: “From the camera’s perspective, estimate how far the closest point of the cabinet (highlighted by a red box) is from the camera in real-world distance, in meters. Select the best answer from the options below: A. 1-2, B. 2-3, C. 3-4, D. 4-5, E. 5-6, F. 6-7”.
The preprocessing steps for outdoor dataset creation are summarized below:
  • To ensure the depth estimation task remains non-trivial, only objects with a minimum distance of 8 meters from the camera were considered. This avoids bias toward near-field predictions and better evaluates model precision in far-range perception.
  • Depth values were discretized into bins, and for each class, we selected between 20 and 60 samples per bin to ensure coverage while avoiding overrepresentation. Bins with fewer than 20 samples were discarded. When bins exceeded 60 samples, selection was sorted by bounding box area to prioritize larger, more reliable objects.
  • Image crops were extracted per object while maintaining the following aspect ratio constraint to preserve visual consistency: the width of the crop must be within the range $\text{height},2\times\text{height}]$. This was enforced using the original image dimensions before padding or resizing.
  • Objects touching any edge of the image were excluded to reduce the likelihood of partial occlusion or clipping.
  • Images were padded vertically as needed to preserve the original aspect ratio.
  • The final outdoor absolute depth set contains approximately 6K samples. For each object class and depth bin, an 80/20 train-test split was applied to maintain consistency in evaluation.
  • The question for each outdoor sample: “Estimate the distance from the camera to the closest part of the cyclist (highlighted by a red box) in meters. Choose the best option: A. 8-9, B. 10-11,…, H. 30-31.”
Figure 21: Examples of Absolute Depth AVA Samples.
Relative Depth xia2024large ; mi2024hierarchical ; zhang2025vision . We curate 11.6K image-object pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes, targeting the task of identifying which of two objects in an image is closer to the camera. Each image contains two distinct objects, each annotated with a bounding box. The model is asked to compare their absolute depth and choose the object that appears closer to the camera (Figure 22).
The preprocessing steps for indoor dataset creation are summarized below:
  • Ambiguous object categories were excluded via a manually curated list of label IDs that often lack clear boundaries or meaningful depth interpretations (e.g., wall, floor, ceiling, etc.).
  • Only object pairs with valid bounding boxes (i.e., non-overlapping, fully inside image boundaries) were considered.
  • To ensure perceptual clarity, candidate object pairs were filtered by requiring an absolute depth difference of at least 0.5 meters between them.
  • For a given object class, only those with at least 10 valid pairings were retained, and a maximum of 30 total pairings per label were allowed.
  • After filtering and sampling, we retained 131 object pairs across 5.7K total questions. Images were padded vertically as needed to preserve their original aspect ratio. These were split into 80% train and 20% test sets while maintaining object label and depth-difference balance.
  • Each question is posed as: “Estimate the real-world distances between the objects in this image. Which object is closer the camera, the sink (highlighted by a red box) or the towel (highlighted by a blue box) to the camera? Choose one option from below: 1. red, 2. blue”.
The preprocessing steps for outdoor dataset creation are summarized below:
  • Object annotations were sourced for the following categories: Car, Van, Pedestrian (merged with Person sitting), and Cyclist.
  • Pairs were formed using both intra-class (e.g., Car vs. Car) and inter-class (e.g., Car vs. Pedestrian) combinations.
  • Pairs were retained only if the depth difference between the two objects was at least 0.5 meters, ensuring a meaningful perceptual gap.
  • To avoid ambiguity and incomplete visual evidence, the following filters were applied:
    • Pairs with occluded objects were excluded.
    • Pairs where either object was touching the image edge were discarded.
    • Only crops satisfying the aspect ratio constraint $\text{height}\leq\text{width}\leq 2\times\text{height}$ were included, ensuring visual consistency.
    • Images were padded vertically as needed to preserve their original aspect ratio.
  • After filtering, approximately 6K valid image-object pairs were retained. For each pair type, an 80/20 train-test split was applied while maintaining distributional balance over object categories and depth separations.
  • Each question is posed as: “Which object is closer to the camera, the van (highlighted in red) or the cyclist (highlighted in blue)? Choose one: 1. red, 2. blue.”
Figure 22: Examples of Relative Depth AVA Samples.
Appendix B Experiment Details
The hyperparameters of Stage-1 vision-language alignment pretraining, Stage-2 visual instruction tuning and Stage-3 AVA-Bench evaluation are show in Table 3.
TinyLLaVa AVA-BENCH
Hyperparameter Stage 1 Stage 2 Stage 3
batch size 16 4 4
grad accum steps 4 8 1
LR 1e-3 2e-5 1e-4
LR schedule cosine decay cosine decay
LR warmup ratio 0.03 0.03
weight decay 0 0
epoch 1 10 (20 for localization AVA)
optimizer AdamW AdamW
DeepSpeed stage 3 3
components finetuned Connector Connector + LLM Connector + LoRA on LLM
sample size 558K 665K
Table 3: Hyperparameters of TinyLLaVa and AVA-Bench Evaluation Stage
B.1 Metric Details
Color Recognition.
We use CIEDE2000 luo2001development to calculate the color differences using the colour Python library. Specifically, we convert the predictions and ground-truths from CIE XYZ tristimulus format to CIE L*a*b* colour space with colour.XYZ_to_Lab, followed by the colour.delta_E(pred, gt, method="CIE 2000") for color differences.
Absolute Depth & Counting.
We use the mean absolute error relative to the ground-truth (see Equation 1). This normalization ensures that errors involving greater distances or counts, which are inherently more challenging, are proportionally penalized less severely.
$\mathrm{MAE/GT}\;=\;\frac{1}{N}\sum_{i=1}^{N}\frac{\bigl{|}\,y_{i}-\hat{y}_{i}\,\bigr{|}}{y_{i}}.$ (1)
where N is the number of test samples, $y_{i}$ is the ground-truth (depth or count) for sample $i$ and $\hat{y}_{i}$ the model prediction.
Localization.
We use the Generalized Intersection-over-Union (GIoU rezatofighi2019generalized , Equation 2):
$\mathrm{GIoU}(A,B)=\frac{|A\cap B|}{|A\cup B|}\;-\;\frac{\bigl{|}\,C\setminus(A\cup B)\bigr{|}}{|C|}\,.$ (2)
where $A$ and $B$ are the prediction and ground-truth bounding boxes and $C$ is the smallest (axis–aligned) enclosing box of $A\cup B$.
OCR.
We evaluate OCR performance with Average Normalized Levenshtein Similarity (ANLS) biten2019scene :
$\mathrm{NLS}(p,g)$ $=1-\frac{\mathrm{Lev}(p,g)}{\max\bigl{(}|p|,|g|\bigr{)}},$ (3)
$\mathrm{ANLS}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathrm{NLS}(p_{i},g_{i})=\frac{1}{N}\sum_{i=1}^{N}\Biggl{(}1-\frac{\mathrm{Lev}(p_{i},g_{i})}{\max\bigl{(}|p_{i}|,|g_{i}|\bigr{)}}\Biggr{)}.$
where $\mathrm{Lev}(p,g)$ is the (Levenshtein) edit distance and $|.|$ denotes string length and N is the number of testing samples.
Others.
All other AVAs employ standard accuracy metrics.
Appendix C More Results and Analyses
C.1 Detailed Overall Results
The results used for plotting Figure 7 and Figure 8 are presented in Table 4 where the best performance for each is AVA is bold and the second best is in italics. For each VFM, the first row is the performance, and the second row is the rank.
AVA Abs. Depth Rel. Depth OCR Counting Localization Object Fine-grained Scene Action Spatial Emotion Orientation Texture Color Average Ranking
[…]
SigLIP-2 0.0843 97.38 81.18 0.225 0.6738 88.19 90.17 72.60 45.33 99.17 58.39 79.2 94.08 12.25 2.4
3 2 1 1 1 1 3 3 2 4 5 4 2 2
AIMv2 0.1008 95.71 62.44 0.254 0.5896 85.04 91.78 72.86 43.61 98.93 60.68 79.71 94.11 19.61 3.9
9 8 3 3 5 4 1 2 3 5 1 3 1 6
SigLIP-1 0.0953 96.58 80.3 0.229 0.6103 87.84 90.94 73.4 45.76 98.66 59.6 78.91 93.66 12.64 3.6
6 7 2 2 4 2 2 1 1 8 4 5 3 3
CLIP 0.08461 97.04 60.44 0.290 0.5787 85.02 86.83 72.32 41.59 99.28 60.42 77.77 93.25 19.85 4.9
4 4 6 7 7 5 4 5 4 3 2 6 5 7
InternVL-2.5 0.08212 97.00 60.88 0.269 0.5850 83.19 72.83 71.63 36.13 99.31 59.99 75.71 91.21 20.09 5.4
2 5 5 5 6 7 7 6 7 2 3 7 7 8
RADIOv2.1 0.07645 97.92 62.44 0.257 0.6617 84.90 85.44 72.50 38.77 99.48 56.59 83.71 93.32 17.16 3.6
1 1 4 4 2 6 6 4 5 1 6 2 4 5
DINOv2 0.08469 97.25 9.97 0.272 0.6598 86.31 85.5 70.99 37.45 98.90 54.44 85.54 93.06 21.54 5.3
5 3 7 6 3 3 5 7 6 6 7 1 6 9
SAM 0.09792 94.13 9.79 0.313 0.5216 76.68 40.06 58.28 17.25 98.76 36.88 69.37 81.02 9.87 7.8
8 9 8 8 8 8 8 9 9 7 9 8 9 1
MiDas-3.0 0.09563 96.63 7.72 0.336 0.4490 75.05 32.83 60.19 18.25 96.86 40.40 67.37 86.38 13.28 8
7 6 9 9 9 9 9 8 8 9 8 9 8 4
Table 4: The detailed overall results where the best performance for each is AVA is bold and the second best is in italics. For each VFM, the first row is the performance, and the second row is the rank. Arrows indicate whether lower ($\downarrow$) or higher ($\uparrow$) is better.
C.2 Detailed Analyses for Each AVA
In the main results, we reported the overall performance of VFMs across various AVAs. However, aggregate metrics can sometimes obscure important nuanced insights. To gain deeper understanding, we conduct detailed analyses by partitioning test samples based on specific criteria (e.g., object size in localization tasks) and examining whether these subgroup trends align with overall performance. Generally, the detailed analyses affirm overall trends, but notable exceptions exist, particularly in localization.
Localization.
We split localization testing samples based on normalized bounding box sizes (relative to image size), where 0.1 indicates an object occupies 10% of the image area. As illustrated in Figure 10, VFMs surprisingly exhibit minimal performance differences when localizing large objects (0.3–0.5). Conversely, performance disparities amplify as object size decreases, revealing significant weaknesses in MiDas and SAM for smaller objects. Consequently, the lower overall performance of MiDas and SAM is predominantly due to poor handling of small targets. Practitioners should thus consider object size distributions when selecting VFMs; SAM and MiDas remain viable if target objects are predominantly large .
Counting.
Counting performance is generally consistent across different datasets and count ranges (Figure 23). A notable exception is SAM, whose accuracy notably improves in denser scenarios.
Figure 23: Detail results for counting for overall and different splits based on datasets and ground-truth count range. Lower MAE/GT is better.
Emotion.
Emotion recognition results exhibit remarkable consistency, with rankings and relative performances highly stable across emotion categories (see Figure 24).
Figure 24: Detail results for emotion for overall and different splits based on emotion types.
Orientation.
Orientation performance remains consistent overall, with some intriguing exceptions. Specifically, VFMs universally achieve near-perfect accuracy for top-view images, presumably due to the distinctive nature of this viewpoint compared to side or frontal views (see Figure 25).
Figure 25: Detail results for orientation for overall and different splits based on viewpoint directions.
Absolute Depth.
Overall, absolute depth performance is stable, though specific VFMs exhibit distinctive trends. SAM notably struggles with near objects but improves significantly with increased distance. Conversely, RADIO demonstrates an opposite pattern, excelling with nearer objects but deteriorating with greater distances (see Figure 26).
Figure 26: Detail results for absolute depth for overall and different splits based on scene type and object count range. Lower MAE/GT is better.
Relative Depth.
Relative depth estimation shows uniformly high performance across VFMs, consistently surpassing 90% accuracy. SAM, however, underperforms notably in interior scenes, consistent with the earlier observation in absolute depth and counting that SAM handles smaller, exterior objects better (see Figure 27).
Figure 27: Detail results for relative depth for overall and different splits based on scene type.
Fine-grained Classification.
Fine-grained classification results are consistently robust across datasets, with the exception of SAM and MiDas, both of which are known to lack semantically rich features espinosa2024there ; chen2023segment , resulting in poorer performance (see Figure 28).
Figure 28: Detail results for fine-grained for overall and different splits based on dataset type.
Scene Recognition.
Scene recognition performance is uniformly consistent across all evaluated datasets, echoing the patterns observed in fine-grained classification, where SAM and MiDas again lag behind other VFMs (see Figure 29).
Figure 29: Detail results for scene for overall and different splits based on datasets.
OCR.
OCR results show consistent patterns with those reported in Section 5.2, highlighting significant underperformance by non-language-aligned VFMs, such as DINOv2 and SAM. Notably, we observe that relative performances across VFMs are stable on short texts (length << 20). However, performance for CLIP and AIM sharply declines with longer text sequences (length >> 20), indicating potential limitations in handling extensive textual information (see Figure 30).
Figure 30: Detail results for counting for overall and different splits based on dataset domain and character length. Higher ANLS is better.
Appendix D Related Works
D.1 VFM Evaluation
Existing evaluation VFM protocols generally fall into two categories. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification han2022survey and COCO for detection or segmentation thisanke2023semantic . For example, DINOv2 oquab2023dinov2 is evaluated on image and video classification, instance recognition, image retrieval, semantic segmentation, and depth estimation. For each task, a task-specific head is trained.
To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks liu2023visual ; zhu2023minigpt ; chowdhery2023palm . For example, in addition to the traditional task-specific evaluation, AIMv2 fini2024multimodal and RADIO  ranzinger2024radio follow the LLM-based evaluation and use a Llama-3.0(8B) grattafiori2024llama and a Vicuna-1.5(7B) zheng2023judging respectively, demonstrating a shift towards generalized multimodal evaluation.
D.2 Atomic Visual Abilities
As discussed in subsection A.1, foundational visual concepts—such as number, color, texture, object identity, and spatial relations—have long been recognized as crucial building blocks in compositional Text-to-Image (T2I) benchmarks. Given their foundational role in generation tasks, these primitives naturally underpin perceptual tasks as well. For example, a concept like ’number’ directly translates into the perceptual task of counting.
A recent work chae24decomposing introduced AVSBench to evaluate whether MLLMs understand basic geometric features, including angle, boundary, orthogonality, and curvature, which they refer to as atomic visual skills. However, AVSBench primarily targets geometric comprehension abilities required for geometric diagrams arising in high-school level mathematics. Moreover, AVSBench provides only test data for MLLMs without addressing potential mismatches between training and test data distributions—an issue highlighted in Section 2.2. Consequently, mispredictions in AVSBench evaluations may result from data distribution mismatches rather than genuine visual deficiencies in VFM. In contrast, AVA-Bench explicitly emphasizes atomic visual abilities essential for general visual reasoning tasks commonly encountered in real-world scenarios. By aligning training and evaluation data distributions, AVA-Bench ensures that evaluation outcomes reliably reflect genuine visual perceptual capabilities of VFMs.
Additionally, a concurrent work wu2025compact defines a set of atomic visual capabilities analogous to ours. However, their goal fundamentally differs from ours: wu2025compact aims to build a visual compositional tuning data recipe that builds complex capabilities from simple atomic capabilities, which can significantly reduce instruction-tuning data volume while maintaining strong performance. In contrast, AVA-Bench’s objective is to systematically evaluate VFMs against atomic visual abilities, pinpointing their exact strengths and weaknesses, and providing a comprehensive diagnostic tool to advance the continual development mai2022online ; lomonaco2022cvpr ; mai2021supervised ; shim2021online ; mai2020batch of robust vision foundation models.